A body of mass 45 kg is moving with a constant velocity of 10m/sec . A constant force acts on the body for 4 seconds and the speed of the body becomes 2m/sec in opposite direction . calculate the acceleration produced.

Grade:12

2 Answers

Arun
6 years ago
Dear student
 
initial velocity(u)= 10m/s
Final velocity(v)= -2m/s
Force=m(v-u)/t= 45(-2-10)/4= -135N
force=m*a
a=f/m= -135/45= -3m /s²
 
Regards
Arun (askIITians forum expert)
Sambhavi saw
2 years ago
The air from a hole in a ballon escapes with uniform rate of 4m/sec. If the ballon contains 10 gram of air and if it shrinks completely in 2.5 sec; then calculate the mean force acting on the ballon.